The lucidabr package^{∗}

Sebastian Rahtz, David Carlisle

1997/03/12

1 Introduction

This file contains L^{A}TEX 2ε package files needed to use Lucida Bright fonts, and .fd files for the fonts as named by Y&Y. .fd and metric files for the Berry-naming scheme are available in fonts/metrics/bh on CTAN.

The Lucida Bright font families: Note that the ‘demi bold’ Lycida fonts are classed as ‘b’ (bold) in L^{A}TEX. The

only ‘bold’ font in the Lucida collection is the bold sans serif font, which is classed as ‘ub’ (ultra bold).

Font File Name  Font Name  L^{A}TEX
Standard Original
hlxb8a  lfd  LucidaFax-Demi  hlx/b/n
hlxbi8a  lfdi  LucidaFax-DemiItalic  hlx/b/it
hlxr8a  lfr  LucidaFax  hlx/m/n
hlxri8a  lfi  LucidaFax-Italic  hlx/m/it
hlhb8a  lbd  LucidaBright-Demi  hlh/b/n
hlhbi8a  lbdi  LucidaBright-DemiItalic  hlh/b/it
hlhr8a  lbr  LucidaBright  hlh/m/n
hlhri8a  lbi  LucidaBright-Italic  hlh/m/it
hlhro8a lbsl  LucidaBrightSlanted  hlh/m/sl
hlhrc8a lbrsc  LucidaBrightSmallcaps  hlh/m/sc
hlhbc8a lbdsc  LucidaBrightSmallcaps-Demi  hlh/b/sc
hlsbi8a  lsdi  LucidaSans-DemiItalic  hls/b/it
hlsb8a  lsd  LucidaSans-Demi  hls/b/n
hlsri8a  lsi  LucidaSans-Italic  hls/m/it
hlsr8a  lsr  LucidaSans  hls/m/n
hlsu8a  lsb  LucidaSans-Bold  hls/ub/n
hlsui8a  lsbi  LucidaSans-BoldItalic  hls/ub/it
hlcrf8a  lbl  LucidaBlackletter  hlcf/m/n
hlcriw8a lbh  LucidaHandwriting-Italic  hlcw/m/n
hlcrie8a lbc  LucidaCalligraphy-Italic  hlce/m/it

∗This file has version number v4.4, last revised 1997/03/12.

Font File Name  Font Name  L^{A}TEX
Standard Original
hlcrn8a lbkr  LucidaCasual  hlcn/m/n
hlcrin8a lbki  LucidaCasual-Italic  hlcn/m/it
hlsrt8a  lstr  LucidaSans-Typewriter  hlst/m/n
hlsrot8a lsto  LucidaSans-TypewriterOblique  hlst/m/sl
hlsbot8a lstbo  LucidaSans-TypewriterBoldOblique hlst/b/sl
hlsbt8a  lstb  LucidaSans-TypewriterBold  hlst/b/n
hlcrt8a  lbtr  LucidaTypewriter  hlct/m/n
hlcbt8a lbtb  LucidaTypewriterBold  hlct/b/n
hlcrot8a lbto  LucidaTypewriterOblique  hlct/m/sl
hlcbot8a lbtbo  LucidaTypewriterBoldOblique  hlct/b/sl
hlcra  lbma  LucidaNewMath-Arrows  hlcm/m/n
hlcba  lbmad  LucidaNewMath-Arrows-Demi  hlcm/b/n
hlcrv  lbme  LucidaNewMath-Extension  hlcv/m/n
hlcry  lbms  LucidaNewMath-Symbol  hlcy/m/n
hlcdy  lbmsd  LucidaNewMath-Symbol-Demi  hlcy/b/n
hlcrim  lbmi  LucidaNewMath-Italic  hlcm/m/itx
hlcrima lbmo  LucidaNewMath-AltItalic  hlcm/m/it
hlcdim  lbmdi  LucidaNewMath-DemiItalic  hlcm/b/itx
hlcdima lbmdo  LucidaNewMath-AltDemiItalic  hlcm/b/it
hlcrm  lbmr  LucidaNewMath-Roman  hlcm/m/n
hlcdm  lbmd  LucidaNewMath-Demibold  hlcm/b/n

2 Packages

2.1 Lucmtime Package

Monotype Times with Lucida Math.

1 ⟨∗lucmtime⟩

2 \def\rmdefault{mntx}
3 \def\sfdefault{cmss}
4 \def\ttdefault{cmtt}
5 \def\Mathdefault{mntluc}
6 \DeclareSymbolFont{letters}{OML}{mntluc}{m}{it}
7 \DeclareSymbolFont{operators}{OT1}{mntx}{m}{n}
8 \SetSymbolFont{letters}{normal}{OML}{mntluc}{m}{it}
9 \SetSymbolFont{letters}{bold}{OML}{mntluc}{b}{it}
10 \SetSymbolFont{operators}{bold}{OT1}{mntx}{b}{n}
11 \SetSymbolFont{operators}{normal}{OT1}{mntx}{m}{n}

12 ⟨/lucmtime⟩

2.2 Lucmin Package

Adobe Minion with Lucida Math.

13 ⟨∗lucmin⟩

14 \def\rmdefault{zmn}
15 \def\sfdefault{zmy}
16 \def\ttdefault{hlct}

17 \renewcommand{\bfdefault}{b}
18 \def\Mathdefault{zmnluc}
19 \DeclareSymbolFont{letters}{OML}{zmnluc}{m}{it}
20 \DeclareSymbolFont{operators}{OT1}{zmn}{m}{n}
21 \SetSymbolFont{letters}{normal}{OML}{zmnluc}{m}{it}
22 \SetSymbolFont{letters}{bold}{OML}{zmnluc}{b}{it}
23 \SetSymbolFont{operators}{bold}{OT1}{zmn}{b}{n}
24 \SetSymbolFont{operators}{normal}{OT1}{zmn}{m}{n}

25 ⟨/lucmin⟩

2.3 Lucidbrb and lucidbry Packages

Compatibility with earlier releases.

26 ⟨∗lucidbrb⟩

27 \RequirePackage[expert,vargreek]{lucidabr}

28 ⟨/lucidbrb⟩ 29 ⟨∗lucidbry⟩

30 \RequirePackage[LY1]{fontenc}
31 \RequirePackage[expert,vargreek]{lucidabr}

32 ⟨/lucidbry⟩

2.4 Lucidbr and lucbmath Packages

Set text and math with Lucida Bright fonts. (Lucbmath package only sets the math fonts.)

33 ⟨∗lucidabright | lucbmath⟩

34 \newif\iflucida@expert
35 \DeclareOption{expert}{\lucida@experttrue}
36 \DeclareOption{noexpert}{\lucida@expertfalse}

Set up the variant text and math sizes which Y&Y suggest for Lucida. The figures for these two options actually come from Frank Mittelbach (oh great one).

The default is to scale, but two options allow you to revert to normal behaviour, or get even smaller.

37 \DeclareOption{nolucidascale}{%
38 \def\DeclareLucidaFontShape#1#2#3#4#5#6{%
39  \DeclareFontShape{#1}{#2}{#3}{#4}{<->#5}{#6}}}
40 \DeclareOption{lucidascale}{%
41 \def\DeclareLucidaFontShape#1#2#3#4#5#6{%
42 \DeclareFontShape{#1}{#2}{#3}{#4}{%
43 <-5.5>s*[1.04]#5%
44 <5.5-6.5>s*[1.02]#5%
45 <6.5-7.5>s*[.99]#5%
46 <7.5-8.5>s*[.97]#5%
47 <8.5-9.5>s*[.96]#5%
48 <9.5-10.5>s*[.95]#5%
49 <10.5-11.5>s*[.94]#5%
50 <11.5-13>s*[.93]#5%
51 <13-15.5>s*[.92]#5%
52 <15.5-18.5>s*[.91]#5%
53 <18.5-22.5>s*[.9]#5%
54 <22.5->s*[.89]#5%
55 }{#6}}}

56 \DeclareOption{lucidasmallscale}{%
57 \def\DeclareLucidaFontShape#1#2#3#4#5#6{%
58 \DeclareFontShape{#1}{#2}{#3}{#4}{%
59 <-5.5>s*[.98]#5%
60 <5.5-6.5>s*[.96]#5%
61 <6.5-7.5>s*[.94]#5%
62 <7.5-8.5>s*[.92]#5%
63 <8.5-9.5>s*[.91]#5%
64 <9.5-10.5>s*[.9]#5%
65 <10.5-11.5>s*[.89]#5%
66 <11.5-13>s*[.88]#5%
67 <13-15.5>s*[.87]#5%
68 <15.5-18.5>s*[.86]#5%
69 <18.5-22.5>s*[.85]#5%
70 <22.5->s*[.84]#5%
71 }{#6}}}

Choose style of letters. Italic3 is not really italic at all, more a roman font with math spacing. Italic2 is not really slanted but a different style of italic, so use an ‘itx’ shape.

72 \DeclareOption{mathitalic1}{\def\letters@shape{it}}
73 \DeclareOption{mathitalic2}{\def\letters@shape{itx}}
74 \DeclareOption{mathitalic3}{\def\letters@shape{n}}

Choose between slanted and upright lowercase Greek.

75 \DeclareOption{slantedgreek}{\def\lcgreek@alphabet{letters}}
76 \DeclareOption{uprightgreek}{\def\lcgreek@alphabet{mathupright}}
   Enable use of \upalpha and \varGamma.
77 \DeclareOption{vargreek}{\let\upalpha\relax\let\varGamma\relax}

Stop the AMS symbol names being declared.

78 \DeclareOption{noamssymbols}{\let\blacksquare\endinput}

Set up the text encoding used in the operators font.

79 \let\operator@encoding\encodingdefault
80 \DeclareOption{OT1}{\def\operator@encoding{OT1}}
81 \DeclareOption{T1}{\def\operator@encoding{T1}}
82 \DeclareOption{LY1}{\def\operator@encoding{LY1}}

Set up the text encodings (not in the lucmath package).

83 ⟨∗lucidabright⟩

84 \renewcommand{\rmdefault}{hlh}
85 \renewcommand{\sfdefault}{hls}
86 \renewcommand{\ttdefault}{hlst}
87 \renewcommand{\bfdefault}{b}
88 \DeclareOption{seriftt}{\def\ttdefault{hlct}}
89 \DeclareOption{fax}{\def\rmdefault{hlx}}
90 \DeclareOption{casual}{\def\rmdefault{hlcn}}
91 \DeclareOption{calligraphic}{%
92 \normalfont
93 \DeclareFontShape\encodingdefault\rmdefault{m}{it}%
94  {<->ssub*hlce/m/it}{}}
95 \DeclareOption{handwriting}{%
96 \normalfont
97 \DeclareFontShape\encodingdefault\rmdefault{m}{it}%

 98  {<->ssub*hlcw/m/it}{}%
 99 \DeclareFontShape\encodingdefault\rmdefault{b}{it}%
100  {<->ssub*hlcw/m/it}{}}

101 ⟨/lucidabright⟩

This package makes a lot of redefinitions. The warnings can be rather annoying so some package options control whether the information is printed to the terminal or log file. More control can be obtained by loading the tracefnt package.

Just show font errors; Warning and info to the log file. The default for this package.

102 \DeclareOption{errorshow}{%
103 \def\@font@info#1{%
104  \GenericInfo{(Font)\@spaces\@spaces\@spaces\space\space}%
105  {LaTeX Font Info: \space\space\space#1}}%
106  \def\@font@warning#1{%
107  \GenericInfo{(Font)\@spaces\@spaces\@spaces\space\space}%
108  {LaTeX Font Warning: #1}}}

The normal L^{A}TEX default, Font Info to the log file and Font Warning to the terminal.

109 \DeclareOption{warningshow}{%
110 \def\@font@info#1{%
111  \GenericInfo{(Font)\@spaces\@spaces\@spaces\space\space}%
112  {LaTeX Font Info: \space\space\space#1}}%
113  \def\@font@warning#1{%
114  \GenericWarning{(Font)\@spaces\@spaces\@spaces\space\space}%
115  {LaTeX Font Warning: #1}}}

On some machines writing all the log info may slow things down so extra option not to log font changes at all.

116 \DeclareOption{nofontinfo}{%
117 \let\@font@info\@gobble
118 \let\@font@warning\@gobble}
119 \ExecuteOptions{noexpert,lucidascale,slantedgreek,mathitalic1,errorshow}
120 \ProcessOptions

121 ⟨/lucidabright | lucbmath⟩

122 ⟨∗lucbmath⟩

New encoding scheme for Math Arrows font

123 \DeclareFontEncoding{LMR}{}{}
124 \DeclareFontSubstitution{LMR}{hlcm}{m}{n}
125 ⟨!luctim⟩ \DeclareSymbolFont{letters}{OML}{hlcm}{m}{\letters@shape}
126 \iflucida@expert
127 \DeclareSymbolFont{mathupright}{OML}{hlcm}{m}{n}
128 \fi
129 \DeclareSymbolFont{symbols}{OMS}{hlcy}{m}{n}
130 \DeclareSymbolFont{largesymbols}{OMX}{hlcv}{m}{n}

The new Expert set for bold math

131 \iflucida@expert
132 ⟨!luctim⟩ \SetSymbolFont{letters}{bold}{OML}{hlcm}{b}{\letters@shape}
133 \SetSymbolFont{mathupright}{bold}{OML}{hlcm}{b}{n}
134 \SetSymbolFont{symbols}{bold}{OMS}{hlcy}{b}{n}
135 \fi

136 % \DeclareSymbolFont{italics}{\encodingdefault}{\rmdefault}{m}{it}
137 \DeclareSymbolFont{arrows}{LMR}{hlcm}{m}{n}
138 \iflucida@expert
139 % \DeclareSymbolFont{boldarrows}{LMR}{hlcm}{b}{n}
140 \SetSymbolFont{arrows}{bold}{LMR}{hlcm}{b}{n}
141 \fi

142 ⟨/lucbmath⟩ 143 ⟨∗lucbmath⟩ 144 ⟨∗!luctim⟩

145 \DeclareSymbolFont{operators}{\operator@encoding}{\rmdefault}{m}{n}
146 \SetSymbolFont{operators}{bold}{\operator@encoding}{\rmdefault}{b}{n}
147 \SetSymbolFont{operators}{normal}{\operator@encoding}{\rmdefault}{m}{n}

Explicitly redeclare all the alphabets just in case, but differentiate between pure Lucida, and the Times mixture, since those have genuine OT1 mimics.

148 \DeclareMathAlphabet\mathbf  \operator@encoding{\rmdefault}{b}{n}
149 \DeclareMathAlphabet\mathrm  \operator@encoding{\rmdefault}{m}{n}
150 \DeclareMathAlphabet\mathsf  \operator@encoding{\sfdefault}{m}{n}
151 \DeclareMathAlphabet\mathit  \operator@encoding{\rmdefault}{m}{it}
152 \DeclareMathAlphabet\mathtt  \operator@encoding{\ttdefault}{m}{n}
153 \DeclareMathAlphabet\mathfrak\operator@encoding{hlcf}{m}{n}
154 \SetMathAlphabet{\mathbf}{bold}{\operator@encoding}{\rmdefault}{b}{n}
155 \SetMathAlphabet{\mathsf}{bold}{\operator@encoding}{\sfdefault}{b}{n}
156 \SetMathAlphabet{\mathrm}{bold}{\operator@encoding}{\rmdefault}{b}{n}
157 \SetMathAlphabet{\mathit}{bold}{\operator@encoding}{\rmdefault}{b}{it}
158 \SetMathAlphabet{\mathtt}{bold}{\operator@encoding}{\ttdefault}{b}{n}

159 ⟨/!luctim⟩ 160 ⟨∗luctim⟩

161 \DeclareMathAlphabet      {\mathbf}{OT1}{\Mathdefault}{b}{n}
162 \DeclareMathAlphabet      {\mathrm}{OT1}{\Mathdefault}{m}{n}
163 \DeclareMathAlphabet      {\mathsf}{OT1}{\sfdefault}{m}{n}
164 \DeclareMathAlphabet      {\mathit}{OT1}{\Mathdefault}{m}{it}
165 \DeclareMathAlphabet      {\mathtt}{OT1}{\ttdefault}{m}{n}
166 \SetMathAlphabet{\mathbf}{bold}{OT1}{\Mathdefault}{b}{n}
167 \SetMathAlphabet{\mathsf}{bold}{OT1}{\sfdefault}{b}{n}
168 \SetMathAlphabet{\mathrm}{bold}{OT1}{\Mathdefault}{b}{n}
169 \SetMathAlphabet{\mathit}{bold}{OT1}{\Mathdefault}{b}{it}
170 \SetMathAlphabet{\mathtt}{bold}{OT1}{\ttdefault}{b}{n}

171 ⟨/luctim⟩

172 \DeclareSymbolFontAlphabet{\mathbb}{arrows}
173 \DeclareSymbolFontAlphabet{\mathscr}{symbols}
174 \iflucida@expert
175 \DeclareSymbolFontAlphabet{\mathup}{mathupright}
176 \fi
177 \DeclareMathAccent\vec  {\mathord}{letters}{"7E}

Symbols taken from the operators font. Need to be careful here as different encodings may have been used.

First check that the AMS have not been redefining \colon. If it does not have this original plain TEX definition, don’t redefine it below.

178 \mathchardef\@tempb"603A
179 \def\@tempa{T1}
180 \ifx\operator@encoding\@tempa
181 \DeclareMathSymbol{!}{\mathclose}{operators}{"21}

182 \DeclareMathSymbol{:}{\mathrel}{operators}{"3A}
183 \DeclareMathSymbol{;}{\mathpunct}{operators}{"3B}
184 \DeclareMathSymbol{?}{\mathclose}{operators}{"3F}
185 \ifx\colon\@tempb
186  \DeclareMathSymbol{\colon}{\mathpunct}{operators}{"3A}
187 \fi
188 \DeclareMathAccent{\acute}{\mathalpha}{operators}{"1}
189 \DeclareMathAccent{\grave}{\mathalpha}{operators}{"0}
190 \DeclareMathAccent{\ddot}{\mathalpha}{operators}{"4}
191 \DeclareMathAccent{\tilde}{\mathalpha}{operators}{"3}
192 \DeclareMathAccent{\bar}{\mathalpha}{operators}{"9}
193 \DeclareMathAccent{\breve}{\mathalpha}{operators}{"8}
194 \DeclareMathAccent{\check}{\mathalpha}{operators}{"7}
195 \DeclareMathAccent{\hat}{\mathalpha}{operators}{"2}
196 \DeclareMathAccent{\dot}{\mathalpha}{operators}{"A}
197 \else
198 \def\@tempa{OT1}
199 \ifx\operator@encoding\@tempa
200 \DeclareMathSymbol{!}{\mathclose}{operators}{"21}
201 \DeclareMathSymbol{:}{\mathrel}{operators}{"3A}
202 \DeclareMathSymbol{;}{\mathpunct}{operators}{"3B}
203 \DeclareMathSymbol{?}{\mathclose}{operators}{"3F}
204 \ifx\colon\@tempb
205  \DeclareMathSymbol{\colon}{\mathpunct}{operators}{"3A}
206 \fi
207 \DeclareMathAccent{\acute}{\mathalpha}{operators}{"13}
208 \DeclareMathAccent{\grave}{\mathalpha}{operators}{"12}
209 \DeclareMathAccent{\ddot}{\mathalpha}{operators}{"7F}
210 \DeclareMathAccent{\tilde}{\mathalpha}{operators}{"7E}
211 \DeclareMathAccent{\bar}{\mathalpha}{operators}{"16}
212 \DeclareMathAccent{\breve}{\mathalpha}{operators}{"15}
213 \DeclareMathAccent{\check}{\mathalpha}{operators}{"14}
214 \DeclareMathAccent{\hat}{\mathalpha}{operators}{"5E}
215 \DeclareMathAccent{\dot}{\mathalpha}{operators}{"5F}
216 \else
217 \def\@tempa{LY1}
218 \ifx\operator@encoding\@tempa
219 \DeclareMathSymbol{!}{\mathclose}{operators}{"21}
220 \DeclareMathSymbol{:}{\mathrel}{operators}{"3A}
221 \DeclareMathSymbol{;}{\mathpunct}{operators}{"3B}
222 \DeclareMathSymbol{?}{\mathclose}{operators}{"3F}
223 \ifx\colon\@tempb
224  \DeclareMathSymbol{\colon}{\mathpunct}{operators}{"3A}
225 \fi
226 \DeclareMathAccent{\acute}{\mathalpha}{operators}{"13}
227 \DeclareMathAccent{\grave}{\mathalpha}{operators}{"12}
228 \DeclareMathAccent{\ddot}{\mathalpha}{operators}{"7F}
229 \DeclareMathAccent{\tilde}{\mathalpha}{operators}{"7E}
230 \DeclareMathAccent{\bar}{\mathalpha}{operators}{"16}
231 \DeclareMathAccent{\breve}{\mathalpha}{operators}{"15}
232 \DeclareMathAccent{\check}{\mathalpha}{operators}{"14}
233 \DeclareMathAccent{\hat}{\mathalpha}{operators}{"5E}
234 \DeclareMathAccent{\vec}{\mathord}{letters}{"7E}

235 \DeclareMathAccent{\dot}{\mathalpha}{operators}{"5}
236 \else
237 \PackageWarningNoLine{lucidabr}
238  {Unknown Operator Encoding!\MessageBreak
239  Math accents may be wrong: assuming OT1 positions}
240 \fi\fi\fi

This section derives mostly from Berthold Horn’s files lcdmacro.tex and amssymblb.tex \begin{matrix}c \\ ◯ \end{matrix}1991, 1992 Y&Y. All Rights Reserved Original from Version 1.2, 1992 June 14; updated ad hoc.

241 \@ifpackageloaded{amsmath}{%

(From M J Downes): it’s possible the factors 1.5, 2, 2.5, 3, 3.5 should be adjusted for Lucida fonts. But that has to be determined by looking at printed tests which I cannot do at the moment. [mjd,24-Jun-1993]

242 \def\biggg{\bBigg@\thr@@}
243 \def\Biggg{\bBigg@{3.5}}
244 }{%
245 \def\big#1{{\hbox{$\left#1\vbox to8.20\p@{}\right.\n@space$}}}
246 \def\Big#1{{\hbox{$\left#1\vbox to10.80\p@{}\right.\n@space$}}}
247 \def\bigg#1{{\hbox{$\left#1\vbox to13.42\p@{}\right.\n@space$}}}
248 \def\Bigg#1{{\hbox{$\left#1\vbox to16.03\p@{}\right.\n@space$}}}
249 \def\biggg#1{{\hbox{$\left#1\vbox to17.72\p@{}\right.\n@space$}}}
250 \def\Biggg#1{{\hbox{$\left#1\vbox to21.25\p@{}\right.\n@space$}}}
251 \def\n@space{\nulldelimiterspace\z@ \m@th}

252 }

Define some extra large sizes — always done using extensible parts

253 \def\bigggl{\mathopen\biggg}
254 \def\bigggr{\mathclose\biggg}
255 \def\Bigggl{\mathopen\Biggg}
256 \def\Bigggr{\mathclose\Biggg}

Following is only really needed if the roman text font is not LucidaBright. Draw the small sizes of ‘[’ and ‘]’ from math italic instead of roman font

257 \DeclareMathSymbol{[}{\mathopen} {letters}{"86}
258 \DeclareMathDelimiter{[}{letters}{"86}{largesymbols}{"02}
259 \DeclareMathSymbol{]}{\mathclose}{letters}{"87}
260 \DeclareMathDelimiter{]}{letters}{"87}{largesymbols}{"03}

Draw the small sizes of ‘(’ and ‘)’ from math italic instead of roman font

261 \DeclareMathSymbol{(}{\mathopen} {letters}{"84}
262 \DeclareMathDelimiter{(}{letters}{"84}{largesymbols}{"00}
263 \DeclareMathSymbol{)}{\mathclose}{letters}{"85}
264 \DeclareMathDelimiter{)}{letters}{"85}{largesymbols}{"01}

Draw ‘=’ and ‘+’ from symbol font instead of roman

265 \DeclareMathSymbol{=}{\mathrel} {symbols}{"83}
266 \DeclareMathSymbol{+}{\mathbin} {symbols}{"82}

Draw small ‘/’ from math italic instead of roman font

267 \DeclareMathSymbol{/}{\mathord} {letters}{"3D}
268 \DeclareMathDelimiter{/}{letters}{"3D}{largesymbols}{"0E}

Make open face brackets accessible, i.e. [[ and ]]

269 \DeclareMathDelimiter{\ldbrack}

270 {\mathopen}{letters}{"82}{largesymbols}{"82}
271 \DeclareMathDelimiter{\rdbrack}
272 {\mathclose}{letters}{"83}{largesymbols}{"83}

Provide access to surface integral signs (linked from text to display size)

273 \DeclareMathSymbol{\surfintop}{\mathop}{largesymbols}{"90}
274 \def\surfint{\surfintop\nolimits}

Make medium size integrals available (NOT linked to display size)

275 \DeclareMathSymbol{\midintop}{\mathop}{largesymbols}{"92}
276 \def\midint{\midintop\nolimits}
277 \DeclareMathSymbol{\midointop}{\mathop}{largesymbols}{"93}
278 \def\midoint{\midointop\nolimits}
279 \DeclareMathSymbol{\midsurfintop}{\mathop}{largesymbols}{"94}
280 \def\midsurfint{\midsurfintop\nolimits}

Extensible integral (use with \bigg, \Bigg, \biggg, \Biggg etc)

281 \DeclareMathDelimiter{\largeint}
282 {\mathop}{largesymbols}{"5A}{largesymbols}{"95}

To close up gaps in special math characters constructed from pieces

283 \def\joinrel{\mathrel{\mkern-4mu}} % \def\joinrel{\mathrel{\mkern-3mu}}

The \mkern-2.5mu undoes the bogus ‘italic correction’ after joiners in LBMA

284 \DeclareMathSymbol{\relbar@}{\mathord}{arrows}{"2D}
285 \def\relbar{\mathrel{\smash\relbar@}\mathrel{\mkern-2.5mu}}
286 \DeclareMathSymbol{\Relbar@}{\mathrel}{arrows}{"3D}
287 \def\Relbar{\Relbar@\mathrel{\mkern-2.5mu}}

The \mkern4mu undoes the overhang at the ends of the joiners (and more)

288 \def\longleftarrow{\leftarrow\relbar\mathrel{\mkern4mu}}
289 \def\longrightarrow{\mathrel{\mkern4mu}\relbar\rightarrow}
290 \def\Longleftarrow{\Leftarrow\Relbar\mathrel{\mkern4mu}}
291 \def\Longrightarrow{\mathrel{\mkern4mu}\Relbar\Rightarrow}

If amsmath is loaded, need to redefine the arrow fill commands as the relative spacing around \relbar and \rightarrow is not what the AMS code expects.

292 \AtBeginDocument{%
293 \@ifpackageloaded{amsmath}{%
294  \def\rightarrowfill@#1{%
295  \m@th\setboxz@h{$#1\relbar$}\ht\z@\z@
296  $#1\mkern4.5mu\mathrel{\copy\z@}%
297  \kern-\wd\z@
298  \cleaders\hbox{$#1\mkern-2mu\box\z@\mkern-2mu$}\hfill%
299  \mkern-4.5mu %
300  \rightarrow$}%
301  \def\leftarrowfill@#1{%
302  \m@th\setboxz@h{$#1\relbar$}\ht\z@\z@
303  $#1\leftarrow
304  \mkern-4.5mu %
305  \cleaders\hbox{$#1\mkern-2mu\copy\z@\mkern-2mu$}\hfill
306  \kern-\wd\z@
307  \mathrel{\box\z@}\mkern4.5mu$}
308  \def\leftrightarrowfill@#1{\m@th\setboxz@h{$#1\relbar$}\ht\z@\z@
309  $#1\leftarrow
310  \mkern-12mu %

311  \cleaders\hbox{$#1\mkern-2mu\box\z@\mkern-2mu$}\hfill
312  \rightarrow$}}%
313  {}}

Some characters that need construction in CM exist complete in math italic or math symbol font.

314 \let\bowtie\undefined
315 \let\models\undefined
316 \let\doteq\undefined
317 \let\cong\undefined
318 \let\angle\undefined
319 \DeclareMathSymbol{\bowtie}{\mathrel}{letters}{"F6}
320 \DeclareMathSymbol{\models}{\mathrel}{symbols}{"EE}
321 \DeclareMathSymbol{\doteq}{\mathrel}{symbols}{"C9}
322 \DeclareMathSymbol{\cong}{\mathrel}{symbols}{"9B}
323 \DeclareMathSymbol{\angle}{\mathord}{symbols}{"8B}

These need undefining so that we can redeclare them.

324 \let\Box\undefined
325 \let\Diamond\undefined
326 \let\leadsto\undefined
327 \let\neq\undefined
328 \let\hookleftarrow\undefined
329 \let\hookrightarrow\undefined
330 \let\mapsto\undefined
331 \let\notin\undefined
332 \let\rightleftharpoons\undefined

Other characters may be found in LucidaNewMath-Arrows (more negated later).

333 \DeclareMathSymbol{\neq}{\mathrel}{arrows}{"94}
334 \DeclareMathSymbol{\rightleftharpoons}{\mathrel}{arrows}{"7A}
335 \DeclareMathSymbol{\leftrightharpoons}{\mathrel}{arrows}{"79}
336 \DeclareMathSymbol{\hookleftarrow}{\mathrel}{arrows}{"3C}
337 \DeclareMathSymbol{\hookrightarrow}{\mathrel}{arrows}{"3E}
338 \DeclareMathSymbol{\mapsto}{\mathrel}{arrows}{"2C}
339 \def\longmapsto{\mapstochar\longrightarrow}

Special L^{A}TEX character definitions (originally from L^{A}TEX symbol font)

340 \let\Join\undefined
341 \let\rhd\undefined
342 \let\lhd\undefined
343 \let\unrhd\undefined
344 \let\unlhd\undefined
345 \DeclareMathSymbol{\Join}{\mathrel}{letters}{"F6}
346 \DeclareMathSymbol{\rhd}{\mathrel}{letters}{"2E}
347 \DeclareMathSymbol{\lhd}{\mathrel}{letters}{"2F}
348 \DeclareMathSymbol{\unlhd}{\mathrel}{symbols}{"F4}
349 \DeclareMathSymbol{\unrhd}{\mathrel}{symbols}{"F5}
350 \DeclareMathSymbol{\Box}{\mathord}{arrows}{"02}
351 \DeclareMathSymbol{\Diamond}{\mathord}{arrows}{"08}
352 \DeclareMathSymbol{\leadsto}{\mathrel}{arrows}{"8E}
353 \DeclareMathSymbol{\leadsfrom}{\mathrel}{arrows}{"8D}
354 \def\mathstrut{\vphantom{f}}

In n-th root, don’t want the ‘n’ to come too close to the radical

355 \def\r@@t#1#2{\setbox\z@\hbox{$\m@th#1\sqrt{#2}$}%

356 \dimen@\ht\z@ \advance\dimen@-\dp\z@
357 \mkern5mu\raise.6\dimen@\copy\rootbox \mkern-7.5mu\box\z@}

Here are some extra definitions of mathematical symbols and operators:

358 \DeclareMathSymbol{\defineequal}{\mathrel}{symbols}{"D6}
359 %\let\notleq\nleq
360 %\let\notgeq\ngeq
361 \DeclareMathSymbol{\notequiv}{\mathrel}{arrows}{"95}
362 %\let\notprec\nprec
363 %\let\notsucc\nsucc
364 \DeclareMathSymbol{\notapprox}{\mathrel}{arrows}{"98}
365 %\let\notpreceq\npreceq
366 %\let\notsucceq\nsucceq
367 \DeclareMathSymbol{\notasymp}{\mathrel}{arrows}{"F3}
368 \DeclareMathSymbol{\notsubset}{\mathrel}{arrows}{"C6}
369 \DeclareMathSymbol{\notsupset}{\mathrel}{arrows}{"C7}
370 \DeclareMathSymbol{\notsim}{\mathrel}{arrows}{"96}
371 \DeclareMathSymbol{\notsubseteq}{\mathrel}{arrows}{"C8}
372 \DeclareMathSymbol{\notsupseteq}{\mathrel}{arrows}{"C9}
373 \DeclareMathSymbol{\notsimeq}{\mathrel}{arrows}{"97}
374 \DeclareMathSymbol{\notsqsubseteq}{\mathrel}{arrows}{"D4}
375 \DeclareMathSymbol{\notsqsupseteq}{\mathrel}{arrows}{"D5}
376 \DeclareMathSymbol{\notcong}{\mathrel}{arrows}{"99}
377 \DeclareMathSymbol{\notin}{\mathrel}{arrows}{"1D}
378 \DeclareMathSymbol{\notni}{\mathrel}{arrows}{"1F}
379 %\let\notvdash\nvdash
380 %\let\notmodels\nvDash
381 %\let\notparallelparallel
382 %\let\noteq\neq
383 %\let\notless\nless
384 %\let\notgreater\ngtr
385 %\let\notmid\nmid
386 \let\Bbb\mathbb

Normal L^{A}TEX draws upper case (upright) Greek from cmr10 — when using the Cork encoding, that isn’t there.

387 \iflucida@expert

If we have the LucidaBright Expert set, we’ll draw them from the upright math font. That way we can get bold math to work on upright upper case Greek.

Why doesn’t this work?

  \documentclass{article}
  \usepackage{lucidabr}
  $\mathbf{\Sigma}$
  \end{document}

The answer lies in the meaning of \mathbf; as fntguide.tex says, it is for alphabetic switching. The straight lucida style says

   \DeclareMathSymbol{\Sigma}{\mathalpha}{largesymbols}{'326}

and the \mathalpha signifies that the \Sigma can change with the alphabet; so this in fact looks for \char'326 in the “mathbf” alphabet when we ask for that. That is defined with

  \SetMathAlphabet{\mathbf}{bold}{\operator@encoding}{\rmdefault}{b}{n}

ie normal text Lucida bold. It all works in CMR because the text fonts have Greek, which is why the symbols are defined as ; in addition, the alphabets like \mathbf explicitly ask for OT1:

 \DeclareMathAlphabet      {\mathbf}{OT1}{cmr}{bx}{n}

so it works in T1 encoding too.

When we get the symbols from other fonts in Lucida, we should no longer classify the fonts as \mathalpha, since the mechanism doesn’t function. So we use \mathord instead, and you only get bold Greek if you change \mathversion. At least it’s consistent.

If, however, we are using the Times mixture, we can keep \mathalpha, as we have the right font layouts around.

388 ⟨∗!luctim⟩

389 \DeclareMathSymbol{\Gamma}{\mathord}{mathupright}{0}
390 \DeclareMathSymbol{\Delta}{\mathord}{mathupright}{1}
391 \DeclareMathSymbol{\Theta}{\mathord}{mathupright}{2}
392 \DeclareMathSymbol{\Lambda}{\mathord}{mathupright}{3}
393 \DeclareMathSymbol{\Xi}{\mathord}{mathupright}{4}
394 \DeclareMathSymbol{\Pi}{\mathord}{mathupright}{5}
395 \DeclareMathSymbol{\Sigma}{\mathord}{mathupright}{6}
396 \DeclareMathSymbol{\Upsilon}{\mathord}{mathupright}{7}
397 \DeclareMathSymbol{\Phi}{\mathord}{mathupright}{8}
398 \DeclareMathSymbol{\Psi}{\mathord}{mathupright}{9}
399 \DeclareMathSymbol{\Omega}{\mathord}{mathupright}{10}
400 \else

It’s in the extension font (largesymbols)

401 \DeclareMathSymbol{\Gamma}{\mathord}{largesymbols}{'320}
402 \DeclareMathSymbol{\Delta}{\mathord}{largesymbols}{'321}
403 \DeclareMathSymbol{\Theta}{\mathord}{largesymbols}{'322}
404 \DeclareMathSymbol{\Lambda}{\mathord}{largesymbols}{'323}
405 \DeclareMathSymbol{\Xi}{\mathord}{largesymbols}{'324}
406 \DeclareMathSymbol{\Pi}{\mathord}{largesymbols}{'325}
407 \DeclareMathSymbol{\Sigma}{\mathord}{largesymbols}{'326}
408 \DeclareMathSymbol{\Upsilon}{\mathord}{largesymbols}{'327}
409 \DeclareMathSymbol{\Phi}{\mathord}{largesymbols}{'330}
410 \DeclareMathSymbol{\Psi}{\mathord}{largesymbols}{'331}
411 \DeclareMathSymbol{\Omega}{\mathord}{largesymbols}{'332}
412 \fi

413 ⟨/!luctim⟩ 414 ⟨∗luctim⟩

415 \DeclareMathSymbol{\Gamma}{\mathalpha}{mathupright}{0}
416 \DeclareMathSymbol{\Delta}{\mathalpha}{mathupright}{1}
417 \DeclareMathSymbol{\Theta}{\mathalpha}{mathupright}{2}
418 \DeclareMathSymbol{\Lambda}{\mathalpha}{mathupright}{3}
419 \DeclareMathSymbol{\Xi}{\mathalpha}{mathupright}{4}
420 \DeclareMathSymbol{\Pi}{\mathalpha}{mathupright}{5}
421 \DeclareMathSymbol{\Sigma}{\mathalpha}{mathupright}{6}
422 \DeclareMathSymbol{\Upsilon}{\mathalpha}{mathupright}{7}
423 \DeclareMathSymbol{\Phi}{\mathalpha}{mathupright}{8}
424 \DeclareMathSymbol{\Psi}{\mathalpha}{mathupright}{9}
425 \DeclareMathSymbol{\Omega}{\mathalpha}{mathupright}{10}
426 \else

It’s in the extension font (largesymbols)

427 \DeclareMathSymbol{\Gamma}{\mathord}{largesymbols}{'320}
428 \DeclareMathSymbol{\Delta}{\mathord}{largesymbols}{'321}
429 \DeclareMathSymbol{\Theta}{\mathord}{largesymbols}{'322}
430 \DeclareMathSymbol{\Lambda}{\mathord}{largesymbols}{'323}
431 \DeclareMathSymbol{\Xi}{\mathord}{largesymbols}{'324}
432 \DeclareMathSymbol{\Pi}{\mathord}{largesymbols}{'325}
433 \DeclareMathSymbol{\Sigma}{\mathord}{largesymbols}{'326}
434 \DeclareMathSymbol{\Upsilon}{\mathord}{largesymbols}{'327}
435 \DeclareMathSymbol{\Phi}{\mathord}{largesymbols}{'330}
436 \DeclareMathSymbol{\Psi}{\mathord}{largesymbols}{'331}
437 \DeclareMathSymbol{\Omega}{\mathord}{largesymbols}{'332}
438 \fi

439 ⟨/luctim⟩

440 \DeclareMathSymbol{\alpha}{\mathord}{\lcgreek@alphabet}{"0B}
441 \DeclareMathSymbol{\beta}{\mathord}{\lcgreek@alphabet}{"0C}
442 \DeclareMathSymbol{\gamma}{\mathord}{\lcgreek@alphabet}{"0D}
443 \DeclareMathSymbol{\delta}{\mathord}{\lcgreek@alphabet}{"0E}
444 \DeclareMathSymbol{\epsilon}{\mathord}{\lcgreek@alphabet}{"0F}
445 \DeclareMathSymbol{\zeta}{\mathord}{\lcgreek@alphabet}{"10}
446 \DeclareMathSymbol{\eta}{\mathord}{\lcgreek@alphabet}{"11}
447 \DeclareMathSymbol{\theta}{\mathord}{\lcgreek@alphabet}{"12}
448 \DeclareMathSymbol{\iota}{\mathord}{\lcgreek@alphabet}{"13}
449 \DeclareMathSymbol{\kappa}{\mathord}{\lcgreek@alphabet}{"14}
450 \DeclareMathSymbol{\lambda}{\mathord}{\lcgreek@alphabet}{"15}
451 \DeclareMathSymbol{\mu}{\mathord}{\lcgreek@alphabet}{"16}
452 \DeclareMathSymbol{\nu}{\mathord}{\lcgreek@alphabet}{"17}
453 \DeclareMathSymbol{\xi}{\mathord}{\lcgreek@alphabet}{"18}
454 \DeclareMathSymbol{\pi}{\mathord}{\lcgreek@alphabet}{"19}
455 \DeclareMathSymbol{\rho}{\mathord}{\lcgreek@alphabet}{"1A}
456 \DeclareMathSymbol{\sigma}{\mathord}{\lcgreek@alphabet}{"1B}
457 \DeclareMathSymbol{\tau}{\mathord}{\lcgreek@alphabet}{"1C}
458 \DeclareMathSymbol{\upsilon}{\mathord}{\lcgreek@alphabet}{"1D}
459 \DeclareMathSymbol{\phi}{\mathord}{\lcgreek@alphabet}{"1E}
460 \DeclareMathSymbol{\chi}{\mathord}{\lcgreek@alphabet}{"1F}
461 \DeclareMathSymbol{\psi}{\mathord}{\lcgreek@alphabet}{"20}
462 \DeclareMathSymbol{\omega}{\mathord}{\lcgreek@alphabet}{"21}
463 \DeclareMathSymbol{\varepsilon}{\mathord}{\lcgreek@alphabet}{"22}
464 \DeclareMathSymbol{\vartheta}{\mathord}{\lcgreek@alphabet}{"23}
465 \DeclareMathSymbol{\varpi}{\mathord}{\lcgreek@alphabet}{"24}
466 \DeclareMathSymbol{\varrho}{\mathord}{\lcgreek@alphabet}{"25}
467 \DeclareMathSymbol{\varsigma}{\mathord}{\lcgreek@alphabet}{"26}
468 \DeclareMathSymbol{\varphi}{\mathord}{\lcgreek@alphabet}{"27}

‘Individual’ Upright lowercase Greek (not currently activated).

469 ⟨∗upalpha⟩

470 \ifx\upalpha\relax
471 \DeclareMathSymbol{\upalpha}{\mathord}{mathupright}{11}
472 \DeclareMathSymbol{\upbeta}{\mathord}{mathupright}{12}
473 \DeclareMathSymbol{\upgamma}{\mathord}{mathupright}{13}
474 \DeclareMathSymbol{\updelta}{\mathord}{mathupright}{14}
475 \DeclareMathSymbol{\upepsilon}{\mathord}{mathupright}{15}
476 \DeclareMathSymbol{\upzeta}{\mathord}{mathupright}{16}
477 \DeclareMathSymbol{\upeta}{\mathord}{mathupright}{17}

478 \DeclareMathSymbol{\uptheta}{\mathord}{mathupright}{18}
479 \DeclareMathSymbol{\upiota}{\mathord}{mathupright}{19}
480 \DeclareMathSymbol{\upkappa}{\mathord}{mathupright}{20}
481 \DeclareMathSymbol{\uplambda}{\mathord}{mathupright}{21}
482 \DeclareMathSymbol{\upmu}{\mathord}{mathupright}{22}
483 \DeclareMathSymbol{\upnu}{\mathord}{mathupright}{23}
484 \DeclareMathSymbol{\upxi}{\mathord}{mathupright}{24}
485 \DeclareMathSymbol{\uppi}{\mathord}{mathupright}{25}
486 \DeclareMathSymbol{\uprho}{\mathord}{mathupright}{26}
487 \DeclareMathSymbol{\upsigma}{\mathord}{mathupright}{27}
488 \DeclareMathSymbol{\uptau}{\mathord}{mathupright}{28}
489 \DeclareMathSymbol{\upupsilon}{\mathord}{mathupright}{29}
490 \DeclareMathSymbol{\upphi}{\mathord}{mathupright}{30}
491 \DeclareMathSymbol{\upchi}{\mathord}{mathupright}{31}
492 \DeclareMathSymbol{\uppsi}{\mathord}{mathupright}{32}
493 \DeclareMathSymbol{\upomega}{\mathord}{mathupright}{33}
494 \DeclareMathSymbol{\upvarepsilon}{\mathord}{mathupright}{34}
495 \fi

496 ⟨/upalpha⟩

Slanted upright Greek.

497 ⟨∗varGamma⟩

498 \ifx\varGamma\relax
499 \DeclareMathSymbol{\varGamma}{\mathord}{letters}{"00}
500 \DeclareMathSymbol{\varDelta}{\mathord}{letters}{"01}
501 \DeclareMathSymbol{\varTheta}{\mathord}{letters}{"02}
502 \DeclareMathSymbol{\varLambda}{\mathord}{letters}{"03}
503 \DeclareMathSymbol{\varXi}{\mathord}{letters}{"04}
504 \DeclareMathSymbol{\varPi}{\mathord}{letters}{"05}
505 \DeclareMathSymbol{\varSigma}{\mathord}{letters}{"06}
506 \DeclareMathSymbol{\varUpsilon}{\mathord}{letters}{"07}
507 \DeclareMathSymbol{\varPhi}{\mathord}{letters}{"08}
508 \DeclareMathSymbol{\varPsi}{\mathord}{letters}{"09}
509 \DeclareMathSymbol{\varOmega}{\mathord}{letters}{"0A}
510 \fi

511 ⟨/varGamma⟩

Definitions for math symbols and operators (normally found in the AMS symbol fonts) using LucidaNewMath fonts MSAM* equivalents:

Stop here if noamssymbols option given.

512 \ifx\blacksquare\endinput\endinput\fi
513 \DeclareMathSymbol{\boxdot}{\mathbin}{symbols}{"ED}
514 \DeclareMathSymbol{\boxplus}{\mathbin}{symbols}{"EA}
515 \DeclareMathSymbol{\boxtimes}{\mathbin}{symbols}{"EC}
516 \DeclareMathSymbol{\square}{\mathord}{arrows}{"02}
517 \DeclareMathSymbol{\blacksquare}{\mathord}{arrows}{"03}
518 \DeclareMathSymbol{\centerdot}{\mathbin}{arrows}{"E1}
519 \DeclareMathSymbol{\lozenge}{\mathord}{arrows}{"08}
520 \DeclareMathSymbol{\blacklozenge}{\mathord}{arrows}{"09}
521 \DeclareMathSymbol{\circlearrowright}{\mathrel}{arrows}{"8C}
522 \DeclareMathSymbol{\circlearrowleft}{\mathrel}{arrows}{"8B}
523 \DeclareMathSymbol{\rightleftharpoons}{\mathrel}{arrows}{"7A}
524 \DeclareMathSymbol{\leftrightharpoons}{\mathrel}{arrows}{"79}
525 \DeclareMathSymbol{\boxminus}{\mathbin}{symbols}{"EB}

526 \DeclareMathSymbol{\Vdash}{\mathrel}{symbols}{"F0}
527 \DeclareMathSymbol{\Vvdash}{\mathrel}{letters}{"D3}
528 \DeclareMathSymbol{\vDash}{\mathrel}{symbols}{"EE}
529 \DeclareMathSymbol{\twoheadrightarrow}{\mathrel}{arrows}{"25}
530 \DeclareMathSymbol{\twoheadleftarrow}{\mathrel}{arrows}{"23}
531 \DeclareMathSymbol{\leftleftarrows}{\mathrel}{arrows}{"71}
532 \DeclareMathSymbol{\rightrightarrows}{\mathrel}{arrows}{"73}
533 \DeclareMathSymbol{\upuparrows}{\mathrel}{arrows}{"72}
534 \DeclareMathSymbol{\downdownarrows}{\mathrel}{arrows}{"74}
535 \DeclareMathSymbol{\upharpoonright}{\mathrel}{arrows}{"75}
536 \DeclareMathSymbol{\downharpoonright}{\mathrel}{arrows}{"77}
537 \DeclareMathSymbol{\upharpoonleft}{\mathrel}{arrows}{"76}
538 \DeclareMathSymbol{\downharpoonleft}{\mathrel}{arrows}{"78}
539 \DeclareMathSymbol{\rightarrowtail}{\mathrel}{arrows}{"29}
540 \DeclareMathSymbol{\leftarrowtail}{\mathrel}{arrows}{"28}
541 \DeclareMathSymbol{\leftrightarrows}{\mathrel}{arrows}{"6E}
542 \DeclareMathSymbol{\rightleftarrows}{\mathrel}{arrows}{"6D}
543 \DeclareMathSymbol{\Lsh}{\mathrel}{arrows}{"7B}
544 \DeclareMathSymbol{\Rsh}{\mathrel}{arrows}{"7D}
545 \DeclareMathSymbol{\rightsquigarrow}{\mathrel}{arrows}{"8E}
546 \DeclareMathSymbol{\leftsquigarrow}{\mathrel}{arrows}{"8D}
547 \DeclareMathSymbol{\leftrightsquigarrow}{\mathrel}{arrows}{"91}
548 \DeclareMathSymbol{\looparrowleft}{\mathrel}{arrows}{"3F}
549 \DeclareMathSymbol{\looparrowright}{\mathrel}{arrows}{"40}
550 \DeclareMathSymbol{\circeq}{\mathrel}{symbols}{"D0}
551 \DeclareMathSymbol{\succsim}{\mathrel}{symbols}{"E1}
552 \DeclareMathSymbol{\gtrsim}{\mathrel}{symbols}{"DD}
553 \DeclareMathSymbol{\gtrapprox}{\mathrel}{letters}{"DB}
554 \DeclareMathSymbol{\multimap}{\mathrel}{letters}{"C7}
555 \DeclareMathSymbol{\image}{\mathrel}{letters}{"C6}
556 \DeclareMathSymbol{\original}{\mathrel}{letters}{"C5}
557 \DeclareMathSymbol{\therefore}{\mathrel}{symbols}{"90}
558 \DeclareMathSymbol{\because}{\mathrel}{symbols}{"91}
559 \DeclareMathSymbol{\doteqdot}{\mathrel}{symbols}{"CA}
560 \DeclareMathSymbol{\triangleq}{\mathrel}{symbols}{"D5}
561 \DeclareMathSymbol{\precsim}{\mathrel}{symbols}{"E0}
562 \DeclareMathSymbol{\lesssim}{\mathrel}{symbols}{"DC}
563 \DeclareMathSymbol{\lessapprox}{\mathrel}{letters}{"DA}
564 \DeclareMathSymbol{\eqslantless}{\mathrel}{letters}{"E2}
565 \DeclareMathSymbol{\eqslantgtr}{\mathrel}{letters}{"E3}
566 \DeclareMathSymbol{\curlyeqprec}{\mathrel}{letters}{"E6}
567 \DeclareMathSymbol{\curlyeqsucc}{\mathrel}{letters}{"E7}
568 \DeclareMathSymbol{\preccurlyeq}{\mathrel}{letters}{"E4}
569 \DeclareMathSymbol{\leqq}{\mathrel}{symbols}{"DA}
570 \DeclareMathSymbol{\leqslant}{\mathrel}{letters}{"E0}
571 \DeclareMathSymbol{\lessgtr}{\mathrel}{symbols}{"DE}
572 \DeclareMathSymbol{\backprime}{\mathord}{letters}{"C8}
573 \DeclareMathSymbol{\axisshort}{\mathord}{arrows}{"39}
574 \DeclareMathSymbol{\risingdotseq}{\mathrel}{symbols}{"CC}
575 \DeclareMathSymbol{\fallingdotseq}{\mathrel}{symbols}{"CB}
576 \DeclareMathSymbol{\succcurlyeq}{\mathrel}{letters}{"E5}
577 \DeclareMathSymbol{\geqq}{\mathrel}{symbols}{"DB}
578 \DeclareMathSymbol{\geqslant}{\mathrel}{letters}{"E1}
579 \DeclareMathSymbol{\gtrless}{\mathrel}{symbols}{"DF}

580 \let\sqsubset\undefined
581 \let\sqsupset\undefined
582 \DeclareMathSymbol{\sqsubset}{\mathrel}{symbols}{"E4}
583 \DeclareMathSymbol{\sqsupset}{\mathrel}{symbols}{"E5}
584 \DeclareMathSymbol{\vartriangleright}{\mathrel}{letters}{"2E}
585 \DeclareMathSymbol{\vartriangleleft}{\mathrel}{letters}{"2F}
586 \DeclareMathSymbol{\trianglerighteq}{\mathrel}{symbols}{"F5}
587 \DeclareMathSymbol{\trianglelefteq}{\mathrel}{symbols}{"F4}
588 \DeclareMathSymbol{\bigstar}{\mathord}{arrows}{"AB}
589 \DeclareMathSymbol{\between}{\mathrel}{letters}{"F2}
590 \DeclareMathSymbol{\blacktriangledown}{\mathord}{arrows}{"07}
591 \DeclareMathSymbol{\blacktriangleright}{\mathrel}{letters}{"F1}
592 \DeclareMathSymbol{\blacktriangleleft}{\mathrel}{letters}{"F0}
593 \DeclareMathSymbol{\arrowaxisright}{\mathord}{arrows}{"37}
594 \DeclareMathSymbol{\arrowaxisleft}{\mathord}{arrows}{"36}
595 \DeclareMathSymbol{\vartriangle}{\mathrel}{arrows}{"04}
596 \DeclareMathSymbol{\blacktriangle}{\mathord}{arrows}{"05}
597 \DeclareMathSymbol{\triangledown}{\mathord}{arrows}{"06}
598 \DeclareMathSymbol{\eqcirc}{\mathrel}{symbols}{"CF}
599 \DeclareMathSymbol{\lesseqgtr}{\mathrel}{letters}{"E8}
600 \DeclareMathSymbol{\gtreqless}{\mathrel}{letters}{"E9}
601 \DeclareMathSymbol{\lesseqqgtr}{\mathrel}{letters}{"EA}
602 \DeclareMathSymbol{\gtreqqless}{\mathrel}{letters}{"EB}
603 \DeclareMathSymbol{\Rrightarrow}{\mathrel}{arrows}{"6C}
604 \DeclareMathSymbol{\Lleftarrow}{\mathrel}{arrows}{"6A}
605 \DeclareMathSymbol{\veebar}{\mathbin}{letters}{"D2}
606 \DeclareMathSymbol{\barwedge}{\mathbin}{symbols}{"F6}
607 \DeclareMathSymbol{\angle}{\mathord}{symbols}{"8B}
608 \DeclareMathSymbol{\measuredangle}{\mathord}{symbols}{"8C}
609 \DeclareMathSymbol{\sphericalangle}{\mathord}{symbols}{"8D}
610 \DeclareMathSymbol{\varpropto}{\mathrel}{symbols}{"2F} % ?
611 \DeclareMathSymbol{\smallsmile}{\mathrel}{letters}{"5E} % ?
612 \DeclareMathSymbol{\smallfrown}{\mathrel}{letters}{"5F} % ?
613 \DeclareMathSymbol{\Subset}{\mathrel}{symbols}{"F8}
614 \DeclareMathSymbol{\Supset}{\mathrel}{symbols}{"F9}
615 \DeclareMathSymbol{\Cup}{\mathbin}{symbols}{"FA}
616 \DeclareMathSymbol{\Cap}{\mathbin}{symbols}{"FB}
617 \DeclareMathSymbol{\curlywedge}{\mathbin}{symbols}{"84}
618 \DeclareMathSymbol{\curlyvee}{\mathbin}{symbols}{"85}
619 \DeclareMathSymbol{\leftthreetimes}{\mathbin}{letters}{"D0}
620 \DeclareMathSymbol{\rightthreetimes}{\mathbin}{letters}{"D1}
621 \DeclareMathSymbol{\subseteqq}{\mathrel}{letters}{"EE}
622 \DeclareMathSymbol{\supseteqq}{\mathrel}{letters}{"EF}
623 \DeclareMathSymbol{\bumpeq}{\mathrel}{symbols}{"C8}
624 \DeclareMathSymbol{\Bumpeq}{\mathrel}{symbols}{"C7}
625 \DeclareMathSymbol{\lll}{\mathrel}{letters}{"DE}
626 \DeclareMathSymbol{\ggg}{\mathrel}{letters}{"DF}
627 \DeclareMathSymbol{\circledS}{\mathord}{letters}{"CA}
628 \DeclareMathSymbol{\pitchfork}{\mathrel}{letters}{"F3}
629 \DeclareMathSymbol{\dotplus}{\mathbin}{symbols}{"89}
630 \DeclareMathSymbol{\backsim}{\mathrel}{letters}{"F8}
631 \DeclareMathSymbol{\backsimeq}{\mathrel}{letters}{"F9}
632 \DeclareMathSymbol{\complement}{\mathord}{letters}{"94}
633 \DeclareMathSymbol{\intercal}{\mathbin}{letters}{"D9}

634 \DeclareMathSymbol{\circledcirc}{\mathbin}{symbols}{"E6}
635 \DeclareMathSymbol{\circledast}{\mathbin}{symbols}{"E7}
636 \DeclareMathSymbol{\circleddash}{\mathbin}{letters}{"CC}

MSBM* equivalents

637 \DeclareMathSymbol{\lvertneqq}{\mathrel}{arrows}{"DE}
638 \DeclareMathSymbol{\gvertneqq}{\mathrel}{arrows}{"DF}
639 \DeclareMathSymbol{\nleq}{\mathrel}{arrows}{"9C}
640 \DeclareMathSymbol{\ngeq}{\mathrel}{arrows}{"9D}
641 \DeclareMathSymbol{\nless}{\mathrel}{arrows}{"9A}
642 \DeclareMathSymbol{\ngtr}{\mathrel}{arrows}{"9B}
643 \DeclareMathSymbol{\nprec}{\mathrel}{arrows}{"E5}
644 \DeclareMathSymbol{\nsucc}{\mathrel}{arrows}{"E6}
645 \DeclareMathSymbol{\lneqq}{\mathrel}{arrows}{"DC}
646 \DeclareMathSymbol{\gneqq}{\mathrel}{arrows}{"DD}
647 \DeclareMathSymbol{\nleqslant}{\mathrel}{arrows}{"D6}
648 \DeclareMathSymbol{\ngeqslant}{\mathrel}{arrows}{"D7}
649 \DeclareMathSymbol{\lneq}{\mathrel}{arrows}{"DA}
650 \DeclareMathSymbol{\gneq}{\mathrel}{arrows}{"DB}
651 \DeclareMathSymbol{\npreceq}{\mathrel}{arrows}{"E7}
652 \DeclareMathSymbol{\nsucceq}{\mathrel}{arrows}{"E8}
653 \DeclareMathSymbol{\precnsim}{\mathrel}{arrows}{"EB}
654 \DeclareMathSymbol{\succnsim}{\mathrel}{arrows}{"EC}
655 \DeclareMathSymbol{\lnsim}{\mathrel}{arrows}{"E0}
656 \DeclareMathSymbol{\gnsim}{\mathrel}{arrows}{"E2}
657 \DeclareMathSymbol{\nleqq}{\mathrel}{arrows}{"D8}
658 \DeclareMathSymbol{\ngeqq}{\mathrel}{arrows}{"D9}
659 \DeclareMathSymbol{\precneqq}{\mathrel}{arrows}{"E9}
660 \DeclareMathSymbol{\succneqq}{\mathrel}{arrows}{"EA}
661 \DeclareMathSymbol{\precnapprox}{\mathrel}{arrows}{"ED}
662 \DeclareMathSymbol{\succnapprox}{\mathrel}{arrows}{"EE}
663 \DeclareMathSymbol{\lnapprox}{\mathrel}{arrows}{"E3}
664 \DeclareMathSymbol{\gnapprox}{\mathrel}{arrows}{"E4}
665 \DeclareMathSymbol{\nsim}{\mathrel}{arrows}{"96}
666 \DeclareMathSymbol{\ncong}{\mathrel}{arrows}{"99}
667 \DeclareMathSymbol{\diagup}{\mathrel}{arrows}{"0B}
668 \DeclareMathSymbol{\diagdown}{\mathrel}{arrows}{"0C}
669 \DeclareMathSymbol{\varsubsetneq}{\mathrel}{arrows}{"D0}
670 \DeclareMathSymbol{\varsupsetneq}{\mathrel}{arrows}{"D1}
671 \DeclareMathSymbol{\nsubseteqq}{\mathrel}{arrows}{"CA}
672 \DeclareMathSymbol{\nsupseteqq}{\mathrel}{arrows}{"CB}
673 \DeclareMathSymbol{\subsetneqq}{\mathrel}{arrows}{"CE}
674 \DeclareMathSymbol{\supsetneqq}{\mathrel}{arrows}{"CF}
675 \DeclareMathSymbol{\varsubsetneqq}{\mathrel}{arrows}{"D2}
676 \DeclareMathSymbol{\varsupsetneqq}{\mathrel}{arrows}{"D3}
677 \DeclareMathSymbol{\subsetneq}{\mathrel}{arrows}{"CC}
678 \DeclareMathSymbol{\supsetneq}{\mathrel}{arrows}{"CD}
679 \DeclareMathSymbol{\nsubseteq}{\mathrel}{arrows}{"C8}
680 \DeclareMathSymbol{\nsupseteq}{\mathrel}{arrows}{"C9}
681 \DeclareMathSymbol{\nparallel}{\mathrel}{arrows}{"F7}
682 \DeclareMathSymbol{\nmid}{\mathrel}{arrows}{"F6}
683 \DeclareMathSymbol{\nshortmid}{\mathrel}{arrows}{"F4}
684 \DeclareMathSymbol{\nshortparallel}{\mathrel}{arrows}{"F5}
685 \DeclareMathSymbol{\nvdash}{\mathrel}{arrows}{"F8}

686 \DeclareMathSymbol{\nVdash}{\mathrel}{arrows}{"FA}
687 \DeclareMathSymbol{\nvDash}{\mathrel}{arrows}{"F9}
688 \DeclareMathSymbol{\nVDash}{\mathrel}{arrows}{"FB}
689 \DeclareMathSymbol{\ntrianglerighteq}{\mathrel}{arrows}{"F2}
690 \DeclareMathSymbol{\ntrianglelefteq}{\mathrel}{arrows}{"F1}
691 \DeclareMathSymbol{\ntriangleleft}{\mathrel}{arrows}{"EF}
692 \DeclareMathSymbol{\ntriangleright}{\mathrel}{arrows}{"F0}
693 \DeclareMathSymbol{\nleftarrow}{\mathrel}{arrows}{"32}
694 \DeclareMathSymbol{\nrightarrow}{\mathrel}{arrows}{"33}
695 \DeclareMathSymbol{\nLeftarrow}{\mathrel}{arrows}{"66}
696 \DeclareMathSymbol{\nRightarrow}{\mathrel}{arrows}{"68}
697 \DeclareMathSymbol{\nLeftrightarrow}{\mathrel}{arrows}{"67}
698 \DeclareMathSymbol{\nleftrightarrow}{\mathrel}{arrows}{"34}
699 \DeclareMathSymbol{\divideontimes}{\mathbin}{letters}{"F7}
700 \DeclareMathSymbol{\varnothing}{\mathord}{letters}{"9C}
701 \DeclareMathSymbol{\nexists}{\mathord}{arrows}{"20}
702 \DeclareMathSymbol{\Finv}{\mathord}{letters}{"90}
703 \DeclareMathSymbol{\Game}{\mathord}{letters}{"91}
704 \let\mho\undefined
705 \DeclareMathSymbol{\mho}{\mathord}{letters}{"92}
706 \DeclareMathSymbol{\simeq}{\mathrel}{symbols}{"27}
707 \DeclareMathSymbol{\eqsim}{\mathrel}{symbols}{"99}
708 \DeclareMathSymbol{\beth}{\mathord}{letters}{"95}
709 \DeclareMathSymbol{\gimel}{\mathord}{letters}{"96}
710 \DeclareMathSymbol{\daleth}{\mathord}{letters}{"97}
711 \DeclareMathSymbol{\lessdot}{\mathrel}{letters}{"DC}
712 \DeclareMathSymbol{\gtrdot}{\mathrel}{letters}{"DD}
713 \DeclareMathSymbol{\ltimes}{\mathbin}{letters}{"CE}
714 \DeclareMathSymbol{\rtimes}{\mathbin}{letters}{"CF}
715 \DeclareMathSymbol{\shortmid}{\mathrel}{letters}{"F4}
716 \DeclareMathSymbol{\shortparallel}{\mathrel}{letters}{"F5}
717 \DeclareMathSymbol{\smallsetminus}{\mathbin}{letters}{"D8} %?
718 \DeclareMathSymbol{\thicksim}{\mathrel}{symbols}{"18} %?
719 \DeclareMathSymbol{\thickapprox}{\mathrel}{symbols}{"19} %?
720 \DeclareMathSymbol{\approxeq}{\mathrel}{symbols}{"9D}
721 \DeclareMathSymbol{\succapprox}{\mathrel}{letters}{"ED}
722 \DeclareMathSymbol{\precapprox}{\mathrel}{letters}{"EC}
723 \DeclareMathSymbol{\curvearrowleft}{\mathrel}{arrows}{"87}
724 \DeclareMathSymbol{\curvearrowright}{\mathrel}{arrows}{"88}
725 \DeclareMathSymbol{\digamma}{\mathord}{letters}{"46} %?
726 \DeclareMathSymbol{\varkappa}{\mathord}{letters}{"9B}
727 \DeclareMathSymbol{\Bbbk}{\mathord}{arrows}{"6B}
728 \DeclareMathSymbol{\hslash}{\mathord}{letters}{"9D}
729 \DeclareMathSymbol{\hbar}{\mathord}{arrows}{"1B}
730 \DeclareMathSymbol{\backepsilon}{\mathrel}{letters}{"FB} %?
731 \DeclareMathSymbol{\dashrightarrow}{\mathord}{arrows}{"3A}
732 \DeclareMathSymbol{\dashleftarrow}{\mathord}{arrows}{"38}
733 \DeclareMathSymbol{\dashuparrow}{\mathord}{arrows}{"39}
734 \DeclareMathSymbol{\dashdownarrow}{\mathord}{arrows}{"3B}
735 \DeclareMathDelimiter\ulcorner{\mathopen}{arrows}{"70}{arrows}{"70}
736 \DeclareMathDelimiter\urcorner{\mathclose}{arrows}{"71}{arrows}{"71}
737 \DeclareMathDelimiter\llcorner{\mathopen}{arrows}{"78}{arrows}{"78}
738 \DeclareMathDelimiter\lrcorner{\mathclose}{arrows}{"79}{arrows}{"79}
739 \edef\checkmark{\noexpand\mathhexbox{\hexnumber@\symarrows}AC}

740 \edef\circledR{\noexpand\mathhexbox{\hexnumber@\symletters}C9}
741 \edef\maltese{\noexpand\mathhexbox{\hexnumber@\symletters}CB}

Changes to default for \Leftrightarrow. I (SPQR) don’t like 22C, so:

742 \let\Leftrightarrow\undefined
743 \DeclareMathSymbol{\Leftrightarrow}{\mathrel}{arrows}{"61}

Override AMS logo, just to ensure we don’t use any CM fonts!

744 %\def\AmS{{\protect\AmSfont
745 %  A\kern-.1667em\lower.5ex\hbox{M}\kern-.125emS}}
746 ⟨lucidabright | lucbmath⟩%\def\AmSfont{\usefont{OMS}{hlcy}{m}{n}}

747 ⟨/lucbmath⟩

2.5 Lucfont test file

A test file for the Lucida fonts.

748 ⟨∗lucfont⟩

749 \documentclass{article}
750 ⟨T1⟩\usepackage[T1]{fontenc}
751 ⟨LY1⟩\usepackage[LY1]{fontenc}
752 \begin{document}
753 \title{All the Lucida text fonts}
754 \author{prepared by Sebastian Rahtz}
755 \date{February 19th 1995}
756 \maketitle
757 \def\test#1#2#3#4#5{%
758 \item[#1/#2/#3]#4 (#5):
759 {\fontfamily{#1}\fontseries{#2}\fontshape{#3}\selectfont
760 Animadversion for a giraffe costs \pounds123. Wa\ss\ ist
761 das f\"ur ein Klopf?
762 We are often na{\"\i}ve vis-\`{a}-vis
763 the d{\ae}monic ph{\oe}nix's official r\^{o}le in fluffy souffl\'{e}s}

764 }

765

766 \begin{description}
767 \test{hlx}{b}{it}{hlxdi8t}{LucidaFax-DemiItalic}
768 \test{hlx}{b}{n}{hlxd8t}{LucidaFax-Demi}
769 \test{hlx}{m}{it}{hlxrir8t}{LucidaFax-Italic}
770 \test{hlx}{m}{n}{hlxr8t}{LucidaFax}

771

772 \test{hlh}{b}{it}{hlcdib8t}{LucidaBright-DemiItalic}
773 \test{hlh}{b}{n}{hlcdb8t}{LucidaBright-Demi}
774 \test{hlh}{m}{it}{hlcrib8t}{LucidaBright-Italic}
775 \test{hlh}{m}{n}{hlcrb8t}{LucidaBright}

776

777 \test{hlce}{m}{it}{hlcrie8t}{LucidaCalligraphy-Italic}

778

779 \test{hlcf}{m}{n}{hlcrf8t}{LucidaBlackletter}

780

781 \test{hlcn}{m}{it}{hlcrin8t}{LucidaCasual-Italic}
782 \test{hlcn}{m}{n}{hlcrn8t}{LucidaCasual}

783

784 \test{hlst}{b}{n}{hlsbt8t}{LucidaSans-TypewriterBold}
785 \test{hlst}{b}{sl}{hlsbot8t}{LucidaSans-TypewriterBoldOblique}

787 \test{hls}{ub}{it}{hlsbi8t}{LucidaSans-BoldItalic}
788 \test{hls}{ub}{n}{hlsb8t}{LucidaSans-Bold}
789 \test{hls}{b}{it}{hlsdi8t}{LucidaSans-DemiItalic}
790 \test{hls}{b}{n}{hlsd8t}{LucidaSans-Demi}
791 \test{hls}{m}{it}{hlsri8t}{LucidaSans-Italic}
792 \test{hls}{m}{n}{hlsr8t}{LucidaSans}

793

794 \test{hlct}{b}{n}{hlcbt8t}{LucidaTypewriterBold}
795 \test{hlct}{b}{sl}{hlcbot8t}{LucidaTypewriterOblique}
796 \test{hlcw}{m}{it}{hlcriw8t}{LucidaHandwriting-Italic}

797

798 \end{description}
799 \end{document}

800 ⟨/lucfont⟩